OpenStudy (chillout):

Solve the inequalities: log2(x+1) <= 3 -x

OpenStudy (fifciol):

in real numbers?

OpenStudy (chillout):

Yes.

OpenStudy (fifciol):

x+1>0 , x>-1 \[\log_2(x+1)\le3 \] \[x+1 \le 2^3\] \[x \le 7\] so \[x \in (-1;7>\] first inequality

OpenStudy (chillout):

First one checks with my answer.

OpenStudy (fifciol):

second \[x^2>-x\]and \[x^2<2x+1\] \[x^2+x>0 \] x(x+1)>0 and x^2-2x-1<0 x^2-2x-1=(x+sqrt2-1)(x-sqrt2-1)   \[x \in ((1-\sqrt2);0)\cup(1;(1+\sqrt2))\]